# Causal-Informed Contrastive Learning: Towards Bias-Resilient Pre-training under Concept Drift

### v1

@@ 1 Introduction
Contrastive learning has proven to be highly effective in pre-training large-scale models, especially in large vision models exemplified by frameworks like SimCLR [1, 2], MoCo series [3, 4], DINO series [5, 6]. However, with the ongoing scaling of large models, data hunger for contrastive learning is raising more attention in the community towards pre-training effectively from drift data. It could be caused by long-tailed data, noise, and domain shift, where concept drift [7, 8] is utilized to uniformly summarize this phenomenon of unpredictable distribution changes in the pre-training through contrastive learning. Hence, a pertinent question emerges: beyond the existing contrastive learning methods, can contrastive paradigm learn from drift pre-training? In this work, we aim to bridge this gap by providing a systematic analysis of the above question. Our findings highlight critical vulnerabilities of the current contrastive pre-training paradigm in adapting to these challenges, underscoring the need for novel strategies to enhance their robustness in drift data streams. More related works are provided in Appendix A.
@@
Current contrastive pre-training methods predominantly adhere to the paradigm of comparing two distinct views of the same object, typically derived from different encoders. Especially in terms of large vision models, student-teacher structures are used to extract different features from two views of images for contrasting agreement, like DINO [5], MoCo v3 [4]. The student network is propelled by the contrastive loss, whereas the teacher network is momentum updated through an exponential moving average (EMA) of the student parameters. In the context of drift pre-training, this contrastive manner will be subtly affected by the concept drift of data streams and be accumulated through momentum, leading to amplified bias. We utilize long-tailed drift as a representative case to demonstrate the impact of concept drift on contrastive pre-training. Illustrated in Fig. 1, we visualize the feature space to investigate how long-tailed drift impacts the contrastive pre-training of MoCo v3 [4]. ImageNet [9] and ImageNet-LT [10] are separately leveraged as balanced and long-tailed sources for contrastive pre-training, where the test data for generating feature space is balanced. Categories are differentiated by varying colors ranging from dark to light, indicating a spectrum from categories with few training samples to those with abundant samples. It is evident from Fig. 1(b), under imbalanced drift pre-training, the tail categories are confined to a small portion of the feature space, whereas the majority is dominated by the head categories with ample samples. The model is steered away from the center of balanced feature space, due to the accumulated bias in momentum update.
@@
Under the above assumption, we formulate the causal relationships [11] among samples $X$, predictions $Y$, the concept drift $D$ within data streams and the bias $B$ in the pre-trained teacher network with momentum update in a causal graph as shown in Fig.2(a). We find the concept drift $D$ within data streams is a confounder of $X$ and $Y$, where the backdoor [12] path $X\leftarrow D\rightarrow B\rightarrow Y$ leads to the false correlation, and the mediation [13] path $X\rightarrow B\rightarrow Y$ confounds the direct contribution of $X\rightarrow Y$. More analyses and details are given in Section 2.1.
Therefore, we propose a novel contrastive learning-based pre-training method with causal intervention [14] to mitigate the bias caused by concept drift in data streams.
@@ 2.1 Causal Views of Drift Pre-training
Concept drift is a statistical phenomenon wherein the distributional characteristics of a target domain undergo arbitrary changes over the course of training [7]. Regarding the drift pre-training, we formula these unpredictable distribution changes with data stream $S_{0,t}=\{s_{0},...,s_{t}\}$, where $s_{i}=(x_{i},z_{i})$ represents a sample with feature vectors $x_{i}$ and contrastive label $z_{i}$, and $t$ denotes the timestamp in the training data stream, corresponding to the iteration step of the pre-training process. In iteration $t$, $S_{0,t}$ follows a certain distribution $F_{0,t}(x,z)$, thus the drift pre-training can be formalized as:
@@
For analyzing contrastive learning under the drift pre-training, we construct a structural causal graph [12, 14] to formula the causal relationship among sample features ($X$), prediction (Y), the contrastive teacher network ($T$) and the drift in the pre-training ($D$) as illustrated in Fig.2, where $A\rightarrow B$ denotes that $A$ is the causer of $B$. The causal graph of $\{X,Y,D,B\}$ presents the following causal connections:
@@
In the constructed causal graph, nodes $D$ and $B$ are identified as the confounder and the mediator [13], respectively. The confounder $D$ influences both the independent variable $X$ and the outcome variable $Y$ from the backdoor path $X\leftarrow D\rightarrow B\rightarrow Y$, thereby distorting the estimated effect of the sample features on the predictions and resulting in the spurious correlation. For example, in the case of long-tailed drift pre-training, tail samples may be incorrectly associated with head categories through this backdoor path, a misalignment attributable to this spurious correlation. Besides, mediator $B$ conveys the drift effect of sample features $X$ on predictions $Y$ via the pathway $X\rightarrow B\rightarrow Y$, weakening the direct impact of $X$ on $Y$.
@@ 2.2 Causal Interventional Contrastive Objective
Drawing from the aforementioned causal relationships, it can be argued that current contrastive pertaining approaches involving the InfoNCE loss [15] essentially leverages the likelihood $P(Y|X)$ to drive the whole network, which is vulnerable by the concept drift confounder $D$ and leads to the emergence of spurious correlations. $P(Y|X)$ can be formulated as:
@@
The de-confound classifier mitigates the confounding influence and captures the genuine causality from $X$ to $Y$, thereby enhancing the quality of the contrastive pre-training. Nevertheless, Eq. (5) necessitates costly sampling to approximate $P(Y|\text{do}(X))$ when implementing it in contrastive pre-training, resulting in impractical training times. Fortunately, the Normalized Weighted Geometric Mean (NWGM) [24, 21, 25] explores an alternative to approximate Eq. (5) in a single feed-forward process as following:
@@ 2.3 Resilient Contrastive Pre-training
Following the student-teacher architecture of MoCo v3 [4], an encoder and a momentum encoder are utilized for causal interventional contrastive pre-training. Within the drift adaption window $[t,t+\tau]$ of the data stream, two distinct crops of each image are utilized under random data augmentation to produce two contrasting views for comparison. The encoder extracts the image feature $v=g(\tilde{x}^{i})$, while the momentum encoder is leveraged for the stability of training as the teacher network, generating the features $k=m(\hat{x}^{i})$. The momentum encoder has the same structure as the encoder, driven by the exponential moving average (EMA) as Eq.(2). Obtained the image feature $v$, an MLP head produce the projection of the encoder feature as $q=h(g(\tilde{x}^{i}))$. Consequently, the intervention module is employed to eliminate latent biases based on features $q$, $v$ and $k$ within the concept drift data stream, and create the contrastive objective $c=\text{Softmax}(q\cdot k)\cdot v$ as Eq. (8). In particular, we use cross-fusion to estimate the concept drift in the data stream, i.e. $q$ and $k$ are obtained from different views, helping the stability of the pre-training. Finally, the InfoNCE loss [15] is adopted to drive the whole network as follows:
@@
where $\tau$ is a temperature hyper-parameter [26]. The sum is over one positive and all negative samples within the drift adaptation window $[t,t+\tau]$ of the data stream.
Intuitively, $q$ akin to a "query" [3], the momentum output $k$ acts as the "key" for sampling key drift of the data stream, and the "value" $v$ resembles a blend of the essential contrastive object and the concept drift. Therefore, the goal of causal intervention is to mitigate the bias of the contrastive objective through the drift sampling within the adaption window. We find the derived causal intervention module has a similar structure to the self-attention mechanism [27]. Thus, it is argued that the causal intervention contrastive learning has the ability to capture long-sequence relationships similar to self-attention, which aids in capturing concept drift within a broader drift adaptation window in contrastive pre-training. However, it is important to highlight that we employ "key" and "query" sampling for concept drift within the data stream to mitigate bias in the "value", which is different from the aim of sequence position modelling in the self-attention mechanism.
@@ 2.4 Simple and Scalable Implementation
Our implementation of the proposed causal interventional contrastive pre-training is highly efficient and simple, requiring minimal specialized operations. As illustrated in the workflow depicted in Fig. 3, we have made only minor adjustments to the MoCo v3 code [4]. These modifications primarily pertain to the interventional steps following the acquisition of image feature embeddings from the encoder and the momentum encoder. Additionally, we have incorporated the masked contrastive strategy proposed in [28] to facilitate a large batch size which enables the acquisition of the border drift adaption window.
Table 1: Evaluation results of fine-tuning on long-tailed classification tasks with ImageNet-LT [10] and iNatualist2018 [29]. The best-performing contrastive pre-training results are highlighted in red. Many, Medium and Few denote the evaluated splits of many-shot ($>$100 training samples), medium-shot (20-100 samples) and few-shot ($<$20 samples). Top-1 accuracy is applied to evaluate the performance of different methods. $\dagger$ denotes methods that are adjusted for the long-tailed fine-tuning.
@@
cRT [30] 61.8 46.2 27.3 49.6 69.0 66.0 63.2 65.2
LWS [30] 60.2 47.2 30.3 49.9 65.0 66.3 65.5 65.9
RIDE [31] 68.2 53.8 36.0 56.9 70.9 72.4 73.1 72.6
PaCo [32] ResNet-50 68.2 58.7 41.0 60.0 70.3 73.2 73.6 73.2
@@
MAE [33] 74.7 48.2 19.4 54.5 79.6 70.8 65.0 69.4
LiVT $\dagger$ [34] ViT-B/16 73.6 56.4 41.0 60.9 78.9 76.5 74.8 76.1
@@
ViT [35] 50.5 23.5 6.9 31.6 65.4 55.3 50.9 54.6
DeiT [36] 70.4 40.9 12.8 48.4 72.9 62.8 55.8 61.0
BYOL [37] 30.6 8.6 0.9 15.9 28.4 22.6 24.5 24.0
DINO [5] 64.8 35.5 11.3 43.3 76.7 67.1 61.2 65.5
DINO v2 [6] 68.4 36.9 11.7 45.4 71.9 62.7 58.2 61.6
MoCo v3 [4] ViT-B/16 70.8 40.7 14.3 48.5 76.6 65.8 62.8 65.6
@@
ImageNet-LT [10]
BYOL [37] 8.5 0.5 0.1 3.5
DINO [5] 47.4 20.5 0.1 27.9
DINO v2 [6] 49.4 23.4 5.7 30.8
MoCo v3 [4] 45.8 18.7 0.0 27.1
@@
iNaturalist2018 [29]
BYOL [37] 7.0 3.1 2.5 3.2
DINO [5] 43.1 33.5 32.3 34.0
DINO v2 [6] 36.3 30.2 30.0 30.7
MoCo v3 [4] 32.9 25.4 22.4 24.9
@@
Furthermore, in comparison to prior approaches, causal contrastive learning presents exceptional scalability. Its scalability is twofold: firstly, it accommodates the inclusion of numerous drift datasets in the pre-training. Following the scaling law [38], augmenting data correlates with an enhancement in model performance. Causal contrastive pre-training guarantees that model performance is bolstered by increasing data without succumbing to the effects of concept drift. Secondly, it strengthens the stability of larger models during drift pre-training, enabling the pre-training of larger models with more parameters and deeper layers.
@@ 3.1 Navigating Tailed Drift Pre-training
We compare our proposed resilient contrastive pre-training method with other models to explicitly demonstrate its superior performance in tailed drift pre-training. Two large-scale datasets, namely the ImageNet-LT [10] and iNaturalist 2018 [29] are utilized as source datasets to perform pre-training, fine-tuning and linear probing, respectively. To effectively illustrate the efficacy of our resilient contrastive pre-training in mitigating tail drift, we follow the criterion and metrics of long-tailed classification [39] to report the Top-1 accuracy across various splits, including Many split with over 100 training samples, Medium split with 20-100 training samples, and Few split with fewer than 20 training samples.
In terms of fine-tuning results shown in Table 1, we compare mainstream contrastive pre-training methods, namely DeiT [36], BYOL [37] , DINO [5] , DINO v2 [6], MoCo v3 [4] with the baseline of ViT [35], and our resilient contrastive learning exhibits superior results beyond other contrastive pre-training methods. Compared to other methods that employ the student-teacher paradigm, such as BYOL, DINO, DINO v2, and MoCo v3, our resilient contrastive pre-training method demonstrates additional performance enhancements on both the Medium and Few splits, especially in the iNaturalist2018 dataset. It is demonstrated that the proposed causal interventional objective significantly alleviates the accumulated bias due to momentum updates of the teacher network. Meanwhile, under the concept drift scenario, BYOL without negative sample pairs has the inferior performance of long-tailed classification to the baseline ViT, while it is not noticeable in equilibrium [4]. We argue that, in drift pre-training, the essence of negative sample pairs is similar to the causal interventional objective, taking into account the causal relationship of all drifts for each positive sample like de-confounding operation in Eq. (5) to delineate the boundary of the feature space. Relying solely on maximizing positive agreement as a pre-training strategy will cause the model to be overwhelmed by the head category, eventually exacerbating concept drift.
@@
Table 3: Evaluation results of domain shift on ImageNet-V2 [40], ImageNet-Sketch [41] and ImageNet-R [42]. The best-performing models are highlighted in red. Many, Medium and Few denote the evaluated splits of many-shot ($>$100 training samples), medium-shot (20-100 samples) and few-shot ($<$20 samples). Top-1 and Top-5 accuracy are applied to evaluate the performance of different methods. We compare our methods with other contrastive pre-training models, including BYOL [37], DINO [5], DINO v2 [6] and MoCo v3 [4].

### v2

# Rolling with the Punches: Resilient Contrastive Pre-training under Non-Stationary Drift

@@ 1 Introduction
Contrastive learning has proven to be highly effective in pre-training large-scale models, especially in large vision models exemplified by frameworks like SimCLR (chenSimpleFrameworkContrastive2020; chenBigSelfSupervisedModels2020), MoCo series (heMomentumContrastUnsupervised2020; chenEmpiricalStudyTraining2021), DINO series (caronEmergingPropertiesSelfSupervised2021; zhouDINOWMWorldModels2024). However, with the ongoing scaling of large models, data hunger for contrastive learning is raising more attention in the community towards pre-training effectively from drift data. It could be caused by long-tailed data, noise, and domain shift, where concept drift (luLearningConceptDrift2019; yang2024adaptingmultimodallargelanguage) is utilized to uniformly summarize this phenomenon of unpredictable distribution changes in the pre-training through contrastive learning. Hence, a pertinent question emerges: beyond the existing contrastive learning methods, can contrastive paradigm learn from drift pre-training? In this work, we aim to bridge this gap by providing a systematic analysis of the above question. Our findings highlight critical vulnerabilities of the current contrastive pre-training paradigm in adapting to these challenges, underscoring the need for novel strategies to enhance their robustness in drift data streams. More related works are provided in Appendix A.
@@
Current contrastive pre-training methods predominantly adhere to the paradigm of comparing two distinct views of the same object, typically derived from different encoders. Especially in terms of large vision models, student-teacher structures are used to extract different features from two views of images for contrasting agreement, like DINO (caronEmergingPropertiesSelfSupervised2021), MoCo v3 (chenEmpiricalStudyTraining2021). The student network is propelled by the contrastive loss, whereas the teacher network is momentum updated through an exponential moving average (EMA) of the student parameters. In the context of drift pre-training, this contrastive manner will be subtly affected by the concept drift of data streams and be accumulated through momentum, leading to amplified bias. We utilize long-tailed drift as a representative case to demonstrate the impact of concept drift on contrastive pre-training. Illustrated in Fig. 1, we visualize the feature space to investigate how long-tailed drift impacts the contrastive pre-training of MoCo v3 (chenEmpiricalStudyTraining2021). ImageNet (russakovskyImagenetLargeScale2015) and ImageNet-LT (liuLargeScaleLongTailedRecognition2019) are separately leveraged as balanced and long-tailed sources for contrastive pre-training, where the test data for generating feature space is balanced. Categories are differentiated by varying colors ranging from dark to light, indicating a spectrum from categories with few training samples to those with abundant samples. It is evident from Fig. 1(b), under imbalanced drift pre-training, the tail categories are confined to a small portion of the feature space, whereas the majority is dominated by the head categories with ample samples. The model is steered away from the center of balanced feature space, due to the accumulated bias in momentum update.
@@
Under the above assumption, we formulate the causal relationships (pearl2014interpretation) among samples $X$, predictions $Y$, the concept drift $D$ within data streams and the bias $B$ in the pre-trained teacher network with momentum update in a causal graph as shown in Fig.2(a). We find the concept drift $D$ within data streams is a confounder of $X$ and $Y$, where the backdoor (pearl1995causal) path $X\leftarrow D\rightarrow B\rightarrow Y$ leads to the false correlation, and the mediation (pearl2022direct) path $X\rightarrow B\rightarrow Y$ confounds the direct contribution of $X\rightarrow Y$. More analyses and details are given in Section 2.1.
Therefore, we propose a novel contrastive learning-based pre-training method with causal intervention (pearl2016causal) to mitigate the bias caused by concept drift in data streams.
@@ 2.1 Causal Views of Drift Pre-training
Concept drift is a statistical phenomenon wherein the distributional characteristics of a target domain undergo arbitrary changes over the course of training (luLearningConceptDrift2019). Regarding the drift pre-training, we formula these unpredictable distribution changes with data stream $S_{0,t}=\{s_{0},...,s_{t}\}$, where $s_{i}=(x_{i},z_{i})$ represents a sample with feature vectors $x_{i}$ and contrastive label $z_{i}$, and $t$ denotes the timestamp in the training data stream, corresponding to the iteration step of the pre-training process. In iteration $t$, $S_{0,t}$ follows a certain distribution $F_{0,t}(x,z)$, thus the drift pre-training can be formalized as:
@@
For analyzing contrastive learning under the drift pre-training, we construct a structural causal graph (pearl1995causal; pearl2016causal) to formula the causal relationship among sample features ($X$), prediction (Y), the contrastive teacher network ($T$) and the drift in the pre-training ($D$) as illustrated in Fig.2, where $A\rightarrow B$ denotes that $A$ is the causer of $B$. The causal graph of $\{X,Y,D,B\}$ presents the following causal connections:
@@
In the constructed causal graph, nodes $D$ and $B$ are identified as the confounder and the mediator (pearl2022direct), respectively. The confounder $D$ influences both the independent variable $X$ and the outcome variable $Y$ from the backdoor path $X\leftarrow D\rightarrow B\rightarrow Y$, thereby distorting the estimated effect of the sample features on the predictions and resulting in the spurious correlation. For example, in the case of long-tailed drift pre-training, tail samples may be incorrectly associated with head categories through this backdoor path, a misalignment attributable to this spurious correlation. Besides, mediator $B$ conveys the drift effect of sample features $X$ on predictions $Y$ via the pathway $X\rightarrow B\rightarrow Y$, weakening the direct impact of $X$ on $Y$.
@@ 2.2 Causal Interventional Contrastive Objective
Drawing from the aforementioned causal relationships, it can be argued that current contrastive pertaining approaches involving the InfoNCE loss (oordRepresentationLearningContrastive2019) essentially leverages the likelihood $P(Y|X)$ to drive the whole network, which is vulnerable by the concept drift confounder $D$ and leads to the emergence of spurious correlations. $P(Y|X)$ can be formulated as:
@@
The de-confound classifier mitigates the confounding influence and captures the genuine causality from $X$ to $Y$, thereby enhancing the quality of the contrastive pre-training. Nevertheless, Eq. (5) necessitates costly sampling to approximate $P(Y|\text{do}(X))$ when implementing it in contrastive pre-training, resulting in impractical training times. Fortunately, the Normalized Weighted Geometric Mean (NWGM) (xuShowAttendTell2016; liuShowDeconfoundTell2022; yangDeconfoundedImageCaptioning2023) explores an alternative to approximate Eq. (5) in a single feed-forward process as following:
@@ 2.3 Resilient Contrastive Pre-training
Following the student-teacher architecture of MoCo v3 (chenEmpiricalStudyTraining2021), an encoder and a momentum encoder are utilized for causal interventional contrastive pre-training. Within the drift adaption window $[t,t+\tau]$ of the data stream, two distinct crops of each image are utilized under random data augmentation to produce two contrasting views for comparison. The encoder extracts the image feature $v=g(\tilde{x}^{i})$, while the momentum encoder is leveraged for the stability of training as the teacher network, generating the features $k=m(\hat{x}^{i})$. The momentum encoder has the same structure as the encoder, driven by the exponential moving average (EMA) as Eq.(2). Obtained the image feature $v$, an MLP head produce the projection of the encoder feature as $q=h(g(\tilde{x}^{i}))$. Consequently, the intervention module is employed to eliminate latent biases based on features $q$, $v$ and $k$ within the concept drift data stream, and create the contrastive objective $c=\text{Softmax}(q\cdot k)\cdot v$ as Eq. (8). In particular, we use cross-fusion to estimate the concept drift in the data stream, i.e. $q$ and $k$ are obtained from different views, helping the stability of the pre-training. Finally, the InfoNCE loss (oordRepresentationLearningContrastive2019) is adopted to drive the whole network as follows:
@@
where $\tau$ is a temperature hyper-parameter (wuUnsupervisedFeatureLearning2018a). The sum is over one positive and all negative samples within the drift adaptation window $[t,t+\tau]$ of the data stream.
Intuitively, $q$ akin to a ”query” (heMomentumContrastUnsupervised2020), the momentum output $k$ acts as the ”key” for sampling key drift of the data stream, and the ”value” $v$ resembles a blend of the essential contrastive object and the concept drift. Therefore, the goal of causal intervention is to mitigate the bias of the contrastive objective through the drift sampling within the adaption window. We find the derived causal intervention module has a similar structure to the self-attention mechanism (vaswaniAttentionAllYou2017). Thus, it is argued that the causal intervention contrastive learning has the ability to capture long-sequence relationships similar to self-attention, which aids in capturing concept drift within a broader drift adaptation window in contrastive pre-training. However, it is important to highlight that we employ ”key” and ”query” sampling for concept drift within the data stream to mitigate bias in the ”value”, which is different from the aim of sequence position modelling in the self-attention mechanism.
@@ 2.4 Simple and Scalable Implementation
Our implementation of the proposed causal interventional contrastive pre-training is highly efficient and simple, requiring minimal specialized operations. As illustrated in the workflow depicted in Fig. 3, we have made only minor adjustments to the MoCo v3 code (chenEmpiricalStudyTraining2021). These modifications primarily pertain to the interventional steps following the acquisition of image feature embeddings from the encoder and the momentum encoder. Additionally, we have incorporated the masked contrastive strategy proposed in (yangMaskedImageContrastive2024) to facilitate a large batch size which enables the acquisition of the border drift adaption window.
Table 1: Evaluation results of fine-tuning on long-tailed classification tasks with ImageNet-LT (liuLargeScaleLongTailedRecognition2019) and iNatualist2018 (vanhornINaturalistSpeciesClassification2018a). The best-performing contrastive pre-training results are highlighted in red. Many, Medium and Few denote the evaluated splits of many-shot ($>$100 training samples), medium-shot (20-100 samples) and few-shot ($<$20 samples). Top-1 accuracy is applied to evaluate the performance of different methods. $\dagger$ denotes methods that are adjusted for the long-tailed fine-tuning.
@@
cRT (kangDecouplingRepresentationClassifier2019) 61.8 46.2 27.3 49.6 69.0 66.0 63.2 65.2
LWS (kangDecouplingRepresentationClassifier2019) 60.2 47.2 30.3 49.9 65.0 66.3 65.5 65.9
RIDE (wangLongtailedRecognitionRouting2020) 68.2 53.8 36.0 56.9 70.9 72.4 73.1 72.6
PaCo (cuiParametricContrastiveLearning2021) ResNet-50 68.2 58.7 41.0 60.0 70.3 73.2 73.6 73.2
@@
MAE (heMaskedAutoencodersAre2022) 74.7 48.2 19.4 54.5 79.6 70.8 65.0 69.4
LiVT $\dagger$ (xuLearningImbalancedData2023) ViT-B/16 73.6 56.4 41.0 60.9 78.9 76.5 74.8 76.1
@@
ViT (dosovitskiyImageWorth16x162021) 50.5 23.5 6.9 31.6 65.4 55.3 50.9 54.6
DeiT (touvronDeiTIIIRevenge2022) 70.4 40.9 12.8 48.4 72.9 62.8 55.8 61.0
BYOL (grillBootstrapYourOwn2020) 30.6 8.6 0.9 15.9 28.4 22.6 24.5 24.0
DINO (caronEmergingPropertiesSelfSupervised2021) 64.8 35.5 11.3 43.3 76.7 67.1 61.2 65.5
DINO v2 (zhouDINOWMWorldModels2024) 68.4 36.9 11.7 45.4 71.9 62.7 58.2 61.6
MoCo v3 (chenEmpiricalStudyTraining2021) ViT-B/16 70.8 40.7 14.3 48.5 76.6 65.8 62.8 65.6
@@
ImageNet-LT (liuLargeScaleLongTailedRecognition2019)
BYOL (grillBootstrapYourOwn2020) 8.5 0.5 0.1 3.5
DINO (caronEmergingPropertiesSelfSupervised2021) 47.4 20.5 0.1 27.9
DINO v2 (zhouDINOWMWorldModels2024) 49.4 23.4 5.7 30.8
MoCo v3 (chenEmpiricalStudyTraining2021) 45.8 18.7 0.0 27.1
@@
iNaturalist2018 (vanhornINaturalistSpeciesClassification2018a)
BYOL (grillBootstrapYourOwn2020) 7.0 3.1 2.5 3.2
DINO (caronEmergingPropertiesSelfSupervised2021) 43.1 33.5 32.3 34.0
DINO v2 (zhouDINOWMWorldModels2024) 36.3 30.2 30.0 30.7
MoCo v3 (chenEmpiricalStudyTraining2021) 32.9 25.4 22.4 24.9
@@
Furthermore, in comparison to prior approaches, causal contrastive learning presents exceptional scalability. Its scalability is twofold: firstly, it accommodates the inclusion of numerous drift datasets in the pre-training. Following the scaling law (muennighoffScalingDataConstrainedLanguage2023), augmenting data correlates with an enhancement in model performance. Causal contrastive pre-training guarantees that model performance is bolstered by increasing data without succumbing to the effects of concept drift. Secondly, it strengthens the stability of larger models during drift pre-training, enabling the pre-training of larger models with more parameters and deeper layers.
@@ 3.1 Navigating Tailed Drift Pre-training
We compare our proposed resilient contrastive pre-training method with other models to explicitly demonstrate its superior performance in tailed drift pre-training. Two large-scale datasets, namely the ImageNet-LT (liuLargeScaleLongTailedRecognition2019) and iNaturalist 2018 (vanhornINaturalistSpeciesClassification2018a) are utilized as source datasets to perform pre-training, fine-tuning and linear probing, respectively. To effectively illustrate the efficacy of our resilient contrastive pre-training in mitigating tail drift, we follow the criterion and metrics of long-tailed classification (yangTdistributedSphericalFeature2023) to report the Top-1 accuracy across various splits, including Many split with over 100 training samples, Medium split with 20-100 training samples, and Few split with fewer than 20 training samples.
In terms of fine-tuning results shown in Table 1, we compare mainstream contrastive pre-training methods, namely DeiT (touvronDeiTIIIRevenge2022), BYOL (grillBootstrapYourOwn2020) , DINO (caronEmergingPropertiesSelfSupervised2021) , DINO v2 (zhouDINOWMWorldModels2024), MoCo v3 (chenEmpiricalStudyTraining2021) with the baseline of ViT (dosovitskiyImageWorth16x162021), and our resilient contrastive learning exhibits superior results beyond other contrastive pre-training methods. Compared to other methods that employ the student-teacher paradigm, such as BYOL, DINO, DINO v2, and MoCo v3, our resilient contrastive pre-training method demonstrates additional performance enhancements on both the Medium and Few splits, especially in the iNaturalist2018 dataset. It is demonstrated that the proposed causal interventional objective significantly alleviates the accumulated bias due to momentum updates of the teacher network. Meanwhile, under the concept drift scenario, BYOL without negative sample pairs has the inferior performance of long-tailed classification to the baseline ViT, while it is not noticeable in equilibrium (chenEmpiricalStudyTraining2021). We argue that, in drift pre-training, the essence of negative sample pairs is similar to the causal interventional objective, taking into account the causal relationship of all drifts for each positive sample like de-confounding operation in Eq. (5) to delineate the boundary of the feature space. Relying solely on maximizing positive agreement as a pre-training strategy will cause the model to be overwhelmed by the head category, eventually exacerbating concept drift.
@@
Table 3: Evaluation results of domain shift on ImageNet-V2 (rechtImageNetClassifiersGeneralize2019), ImageNet-Sketch (wangLearningRobustGlobal2019) and ImageNet-R (hendrycksManyFacesRobustness2021). The best-performing models are highlighted in red. Many, Medium and Few denote the evaluated splits of many-shot ($>$100 training samples), medium-shot (20-100 samples) and few-shot ($<$20 samples). Top-1 and Top-5 accuracy are applied to evaluate the performance of different methods. We compare our methods with other contrastive pre-training models, including BYOL (grillBootstrapYourOwn2020), DINO (caronEmergingPropertiesSelfSupervised2021), DINO v2 (zhouDINOWMWorldModels2024) and MoCo v3 (chenEmpiricalStudyTraining2021).